# Transversal Hamilton cycles in digraph collections

### v1

@@ 5 Extremal case
If $|\tilde{\mathcal{C}}|=0$, then $n$ is odd. In view of Claim 5.7 (i), we know $D_{i}[B\setminus(X^{\prime}\cup V(P^{1}))]=\emptyset$ for all $i\in\mathcal{C}\setminus{\rm col}(P^{1}-A)$. Notice that $|B\setminus(X^{\prime}\cup V(P^{1}))|\geq\frac{n}{2}-28\delta n$ and $|{\rm col}(P^{1}-A)|\leq 27\delta n$. Together with Lemma 5.14 (see the final of this subsection), we know that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
If $|\tilde{\mathcal{C}}|\geq 1$, then based on Claim 5.7 (ii) and the fact that $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, we obtain $|A|,|B|\geq\frac{n}{2}+1$, a contradiction.
@@
In this case, each vertex in $Y\setminus(X\cup X_{Y}^{1})$ belongs to $Y_{i}$ for at least $(1-13\sqrt{\delta})|\mathcal{C}_{1}|$ colors $i\in\mathcal{C}_{1}$. By a similar discussion as in Case 1, we know that there are $|V_{\rm bad}|$ disjoint rainbow directed $P_{3}$ with centers in $V_{\rm bad}$ and both endpoints in the same part. We now consider vertices in $X^{\prime}$. Based on the definition of $X^{\prime}$, we move vertices in $X^{\prime}$ to $Y$ if they are covered by a rainbow directed $P_{3}$ from $Z$ to $Y$ (in Step 2 of Case 1). In order to connect those rainbow directed $P_{3}$ with centers in $X^{\prime}$ into a single rainbow directed path, we balance the number of rainbow directed paths from $A$ to $B$ and those from $B$ to $A$.
@@
There exists a rainbow directed path $P^{0}$ such that its endpoints are not in $V_{\rm bad}\cup X^{\prime}$, $X^{\prime}\subseteq V(P^{0})$ and ${\rm col}(P^{0})\subseteq\mathcal{C}_{1}$, and its length is at most $8\sqrt{\epsilon}n$. Furthermore, such $P^{0}$ can be chosen starting at $A$ and ending at $B$ or two endpoints in the same part.
@@
If $|X^{\prime}|=0$, then by the minimum semi-degree condition, our result holds obviously. Without loss of generality, assume that $|X^{\prime}|\geq 1$ and $|A\cap X^{\prime}|-|B\cap X^{\prime}|\geq 0$. We first select all disjoint non-trivial maximum rainbow directed paths, say $Q_{1},\ldots,Q_{t}$, inside $\{D_{i}[A\cap X^{\prime}]:i\in\mathcal{C}_{1}\}$. Let $t^{\prime}:=|V(Q_{1})|+\cdots+|V(Q_{t})|$ and $Q_{i}:=x_{i}\ldots y_{i}$ for each $i\in[t]$. Next, choose a maximum rainbow directed matching, denoted by $\tilde{M}$, inside $\{D_{i}[A,B]:i\in\mathcal{C}_{1}\setminus{\rm col}(Q_{1}\cup Q_{2}\cup\ldots\cup Q_{t})\}$ avoiding vertices in $\bigcup_{i\in[t]}V(Q_{i}-y_{i})$. Partition $\tilde{M}$ to $M_{1}\cup M_{2}$ such that $V(M_{2})\cap A=V(\tilde{M})\cap A\cap X^{\prime}$. Hence, each vertex in $V(M_{2})\cap A$ or each $Q_{i}$ with $y_{i}\in V(M_{2})\cap A$ can be covered by a rainbow directed path with length at most $|X^{\prime}|+2$ and both endpoints in $B\setminus(V_{\rm bad}\cup X^{\prime})$. Denote $\hat{X}:=(A\cap X^{\prime})\setminus V\left(\bigcup_{i\in[t]}(Q_{i}-y_{i})\cup\tilde{M}\right)$.
If $|\hat{X}|=0$ and $|B\cap X^{\prime}|\geq 1$, then the desired rainbow directed path can be obtained immediately by Claim 5.3 (iii) (note that if there is a rainbow directed path with some endpoint in $V_{\rm bad}$, then it can be extended by at most two vertices so that both of its endpoints remain in the corresponding part and are no longer in $V_{\rm bad}$).
Next, we consider $|\hat{X}|\geq 1$. We are to prove that $|E(M_{1})|\geq|\hat{X}|+1$. Suppose that $|E(M_{1})|\leq|\hat{X}|\leq\sqrt{\epsilon}n$. Together with the definition of $X^{\prime}$, one has $B\cap X^{\prime}\subseteq V(\tilde{M})\cap B$ and $\tilde{M}$ can be chosen such that each matching edge incident to a vertex in $B\cap X^{\prime}$ is not incident to vertices in $A\cap X^{\prime}$.
Choose $u_{1}\in\hat{X}$, $u_{2}\in B\setminus V(\tilde{M})$ and two colors $c_{1},c_{2}\in\mathcal{C}_{1}\setminus{\rm col}(\tilde{M}\cup Q_{1}\cup\ldots\cup Q_{t})$. We consider $N_{c_{1}}^{+}(u_{1})$ and $N_{c_{2}}^{-}(u_{2})$ respectively. It is easy to see that $u_{1}$ has at most $t^{\prime}-t$ out-neighbors inside $A\cap X^{\prime}$ and $N_{c_{1}}^{+}(u_{1},B)=N_{c_{1}}^{+}(u_{1},B\cap V(\tilde{M}))$. Hence
$\left\lceil\frac{n}{2}\right\rceil\leq|N_{c_{1}}^{+}(u_{1})|\leq|A\setminus X^{\prime}|+(t^{\prime}-t)+|N_{c_{1}}^{+}(u_{1},B\cap V(\tilde{M}))|.$ (2)
@@
Notice that, if there exists an edge $w_{1}w_{2}\in E(\tilde{M})$ such that $u_{1}w_{1}\in E(D_{c_{1}})$ and $w_{2}u_{2}\in E(D_{c_{2}})$, then we will find a large rainbow matching, a contradiction. Hence $|N_{c_{1}}^{+}(u_{1},B\cap V(\tilde{M}))|+|N_{c_{2}}^{-}(u_{2},A\cap V(\tilde{M}))|\leq|E(\tilde{M})|$. It follows that $|E(\tilde{M})|\geq|\hat{X}|+|E(M_{2})|+1$, which implies $|E(M_{1})|\geq|\hat{X}|+1$. Thus, together with Claim 5.3 (iii), the desired rainbow directed path $P^{0}$ can be obtained.
Finally, we consider $|\hat{X}|=0$ and $|B\cap X^{\prime}|=0$. If $|E(M_{1})|\geq 1$, then we are done. Otherwise, by the maximality of $\tilde{M}$ and the definition of $X^{\prime}$, we deduce that $D_{i}[A\setminus X^{\prime},B]=\emptyset$. Using a similar argument as (2) and (3), there is a vertex in $A\setminus(V_{\rm bad}\cup X^{\prime})$ that is adjacent to some vertex in $\{x_{1},\ldots,x_{t}\}\cup((A\cap X^{\prime})\setminus V(Q_{1}\cup\ldots\cup Q_{t}))$. Together with Claim 5.3 (iii), the desired rainbow directed path can be obtained immediately. ∎
By a similar discussion as Claim 5.4, we get the following claim.
Claim 5.9.
Assume $X\setminus X^{\prime}=\{x_{1},\ldots,x_{s}\}$ with $s\leq\sqrt{\epsilon}n$. If $x_{i}\in(X\setminus X^{\prime})\cap Y$, then there exist $c_{i}^{1},c_{i}^{2},c_{i}^{3},c_{i}^{4}\in\mathcal{C}_{1}\setminus({\rm col}(P^{0})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i-1]\})$ and $x_{i}^{1},\,x_{i}^{2},\,x_{i}^{3},\,x_{i}^{4}\in Y\setminus(V(P^{0})\cup V_{\rm bad}\cup\{x_{\ell}^{1},\,x_{\ell}^{2},\,x_{\ell}^{3},\,x_{\ell}^{4}:{\ell}\in[i-1]\})$ such that ${x_{i}^{k}x_{i}}\in E(D_{c_{i}^{k}})$ for $k\in[2]$ and ${x_{i}x_{i}^{k}}\in E(D_{c_{i}^{k}})$ for $k\in\{3,4\}$.
By Claim 5.9 and a similar discussion as Steps 1-3 in Case 1, the following hold.
@@
    In $\{D_{i}[A\setminus V(P^{0})]\cup D_{i}[B\setminus V(P^{0})]:i\in\mathcal{C}_{1}\setminus{\rm col}(P^{0})\}$, there are $|V_{\rm bad}\setminus V(P^{0})|$ disjoint rainbow directed $P_{3}$ with centers in $V_{\rm bad}\setminus V(P^{0})$ and both endpoints in the same part. Let $\mathbf{P}$ be a set consisting of those rainbow directed $P_{3}$.
@@
[…]
@@
    Let $q:=|\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime}|$. In the digraph collection $\{D_{i}^{\pm}[A,B]:i\in\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime})\}$, there exists a rainbow directed path $Q$ of length $q-1$ or $q$ that avoids vertices in $\mathbf{P}\cup P^{0}\cup M$.
@@
Next, we give the following claim.
@@
$\geq d_{i}^{+}(z,A_{i})-|A_{1}\Delta A_{i}|-|X_{A}\setminus X_{A}^{1}|-|X^{\prime}|-|V(Q_{1}\cup Q_{2})|-|I_{3-j}|$
@@
Similarly, we have $|N_{{i}}^{-}(z,I_{j})|\geq\frac{n}{4}-5\eta n$. It implies that
@@
Hence, for each $i\in\mathcal{\mathcal{C}}_{a}\cup\{i_{1},i_{2}\}$, one may view $D_{i}^{\pm}[I_{1},I_{2}]$ as an undirected graph $G_{i}[I_{1},I_{2}]$ with minimum degree at least $\frac{n}{4}-10\eta n$, where two vertices are adjacent in $G_{i}[I_{1},I_{2}]$ if and only if there are two opposite arcs between them. By Claim 5.5, there exists a transversal undirected path $Q_{a}$ inside $\mathcal{D}_{a}$ with two endpoints $x_{a}\in N_{G_{i_{1}}}(z_{1},I_{1})$ and $y_{a}\in N_{G_{i_{2}}}(z_{1}^{\prime},I)$ (here $y_{a}\in I_{2}$ if and only if $|I_{1}|=|I_{2}|$).
@@
Using Claim 5.3 (iii), we connect $P^{0}$ (with endpoints in the same part), all rainbow directed paths in $\mathbf{P},M$ and $Q$ (from $A$ to $B$ with length $q-1$) into a single rainbow directed path $P$, whose length is at most $(8\sqrt{\epsilon}+3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$ and endpoints are $x\in A\setminus(V_{\rm bad}\cup X^{\prime})$, $y\in B\setminus(V_{\rm bad}\cup X^{\prime})$. Next, choose an edge $u_{q}v_{q}$ from $D_{c}[B\setminus V(P),A\setminus V(P)]$, where $c\in(\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus{\rm col}(P)$. Consequently, $P$ and $u_{q}v_{q}$ are two disjoint rainbow directed paths satisfying all conditions in Claim 5.10. Thus, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
Using Claim 5.3 (iii), we connect all rainbow paths in $\mathbf{P},M$ and $Q$ (from $B$ to $A$ with length $q$) into a single rainbow directed path $P$, whose length is at most $(3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$ and endpoints are $x\in B\setminus V_{\rm bad}$, $y\in A\setminus V_{\rm bad}$. Consequently, $P$ and $P^{0}$ (from $A$ to $B$) are two disjoint rainbow paths satisfying all conditions in Claim 5.10. Thus, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
If $|X^{\prime}|\geq 2$, then choose $u\in A\cap X^{\prime}$ if $|A\cap X^{\prime}|>|B\cap X^{\prime}|$, and choose $u\in B\cap X^{\prime}$ if $|A\cap X^{\prime}|=|B\cap X^{\prime}|$. Delete $u$ from $X^{\prime}$ when constructing $P^{0}$. Based on the definition of $X^{\prime}$, there exists a rainbow directed $P_{3}$ from $B$ to $A$ with center $u$ that avoids vertices in $P^{0}$. Applying Claim 5.3 (iii) again to connect $P^{0}$ (from $A$ to $B$), all rainbow directed paths in $\mathbf{P},M$ into a single rainbow path $P$, whose length is at most $(8\sqrt{\epsilon}+3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$, with endpoints in different parts. Thus, $P$ and $P_{x}$ are two disjoint rainbow paths satisfying all conditions in Claim 5.10. Therefore, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
Therefore, there exists a set of disjoint rainbow directed path $\mathbf{P}^{0}=\{v_{i}^{1}v_{i}v_{i}^{2}:i\in[t]\}$ with ${\rm col}(v_{i}^{1}v_{i}v_{i}^{2})=\{c_{4i-3},c_{4i-2}\}$ for each $i\in[t]$.
@@
Based on the definition of $X^{\prime}$, move vertices in $X^{\prime}$ to $Y$ if they are covered by a rainbow directed $P_{3}$ from $Z$ to $Y$ (in Step 2 of Case 1). Thus, by avoiding all vertices and colors in $\mathbf{P}^{0}$, there exists a set of disjoint rainbow directed paths $\mathbf{P^{1}}=\{v^{1}vv^{2}:v\in X^{\prime}\}$ with $v^{1}\in Z\setminus(V_{\rm bad}\cup X^{\prime})$ and $v^{2}\in Y\setminus(V_{\rm bad}\cup X^{\prime})$ if $v\in Y\cap X^{\prime}$. Furthermore, at least one of the colors of ${v^{1}v}$ and ${vv^{2}}$ can be chosen from both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. Without loss of generality, assume that $|A|\leq|B|$.
@@
Connect all rainbow directed paths in $\mathbf{P}^{0}\cup\mathbf{P}^{1}$ by using colors in $\mathcal{C}_{1}\cup\mathcal{C}_{2}$ and Claim 5.3 (i)-(iii), one may obtain a rainbow directed path $P^{1}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{1}\in B\setminus V_{\rm bad}$, whose length is at most $4t+4|X^{\prime}|\leq 2\sqrt{\delta}n+4|X^{\prime}|$.
By a similar discussion as Claim 5.4, we obtain the following claim.
Claim 5.11.
Assume $X\setminus X^{\prime}=\{x_{1},\ldots,x_{s}\}$ with $s\leq\sqrt{\epsilon}n$. Then for each $i\in[s]$, there exist $c_{i}^{1},c_{i}^{2},c_{i}^{3},c_{i}^{4}\in(\mathcal{C}_{1}\cup\mathcal{C}_{2})\setminus({\rm col}(P^{1})\cup\{c_{\ell}^{1},c_{\ell}^{2},c_{\ell}^{3},c_{\ell}^{4}:{\ell}\in[i-1]\})$ and $x_{i}^{1},\,x_{i}^{2},\,x_{i}^{3},\,x_{i}^{4}\in V\setminus(V(P^{1})\cup\{x_{\ell}^{1},\,x_{\ell}^{2},\,x_{\ell}^{3},\,x_{\ell}^{4}:{\ell}\in[i-1]\})$ such that: if $x_{i}\in Y$, then ${x_{i}^{j}x_{i}}\in E(D_{c_{i}^{j}})$ for all $j\in[2]$, and ${x_{i}x_{i}^{j}}\in E(D_{c_{i}^{j}})$ for all $j\in\{3,4\}$, and either $x_{i}^{j}\in Y$ and $c_{i}^{j}\in\mathcal{C}_{1}$ for all $j\in[4]$, or $x_{i}^{j}\in Z$ and $c_{i}^{j}\in\mathcal{C}_{2}$ for all $j\in[4]$.
@@
By Claim 5.11, there exists a set of disjoint rainbow directed path $\mathbf{P}^{2}=\{x_{i}^{1}x_{i}x_{i}^{2}:i\in[s]\}$, each of which has colors $c_{i}^{1},c_{i}^{2}$ for $i\in[s]$. Connect all these rainbow directed paths and $P^{1}$ using Claim 5.3 (i)-(ii), and we obtain a rainbow directed path $P^{2}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{2}\in B\setminus V_{\rm bad}$, whose length is at most $|E(P^{1})|+4|X\setminus X^{\prime}|+2\leq(2\sqrt{\delta}+4\sqrt{\epsilon})n+2$.
[…]

### v2

@@ 1. $(\dagger)$
If $|\tilde{\mathcal{C}}|=0$, then $n$ is odd. In view of Claim 6.7 (i), we know $D_{i}[B\setminus(X^{\prime}\cup V(P^{1}))]=\emptyset$ for all $i\in\mathcal{C}\setminus{\rm col}(P^{1}-A)$. Notice that $|B\setminus(X^{\prime}\cup V(P^{1}))|\geq\frac{n}{2}-28\delta n$ and $|{\rm col}(P^{1}-A)|\leq 27\delta n$. Together with Lemma 6.5 (see the end of this subsection), we know that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
If $|\tilde{\mathcal{C}}|\geq 1$, then based on Claim 6.7 (ii) and the fact that $\delta^{0}(\mathcal{D})\geq\frac{n}{2}$, we obtain $|A|,|B|\geq\frac{n+1}{2}$, a contradiction.
@@
In this case, each vertex in $Y\setminus(X\cup X_{Y}^{1})$ belongs to $Y_{i}$ for at least $(1-13\sqrt{\delta})|\mathcal{C}_{1}|$ colors $i\in\mathcal{C}_{1}$. Following the argument from Step 2 of Case 1, we know that there are $|V_{\rm bad}|$ disjoint rainbow directed $P_{3}$ copies, each having its center in $V_{\rm bad}$ and both endpoints in the same part. We now consider vertices in $X^{\prime}$. Based on the definition of $X^{\prime}$, we move a vertex in $X^{\prime}$ to $Y$ whenever it is covered by a rainbow directed $P_{3}$ from $Z$ to $Y$ (in Step 2 of Case 1). In order to connect those rainbow directed $P_{3}$ with centers in $X^{\prime}$ into a single rainbow directed path, we balance the number of rainbow directed paths from $A$ to $B$ and those from $B$ to $A$.
@@
There exists a rainbow directed path $P^{0}$ with length at most $8\sqrt{\epsilon}n$ such that $X^{\prime}\subseteq V(P^{0})$, ${\rm col}(P^{0})\subseteq\mathcal{C}_{1}$, and its endpoints lie outside $V_{\rm bad}\cup X^{\prime}$. Furthermore, $P^{0}$ can be chosen to either start at $A$ and end at $B$, or have both endpoints in the same part.
@@
[…]
Assume that $|\hat{X}|\geq 1$. We are to prove that $|E(M_{1})|\geq|\hat{X}|+1$. Suppose for contradiction that $|E(M_{1})|\leq|\hat{X}|\leq\sqrt{\epsilon}n$. Together with the definition of $X^{\prime}$, one has $B\cap X^{\prime}\subseteq V(\tilde{M})\cap B$ and $\tilde{M}$ can be chosen such that each matching edge incident to a vertex in $B\cap X^{\prime}$ is not incident to vertices in $A\cap X^{\prime}$.
[…]
@@
[…]
Now, we consider $|\hat{X}|=0$ and $|B\cap X^{\prime}|=0$. If $|E(M_{1})|\geq 1$, then we are done. Otherwise, by the maximality of $\tilde{M}$ and the definition of $X^{\prime}$, we deduce that $D_{i}[A\setminus X^{\prime},B]=\emptyset$. Using a similar argument as (3) and (4), there is a vertex in $A\setminus(V_{\rm bad}\cup X^{\prime})$ that is adjacent to some vertex in $\{x_{1},\ldots,x_{t}\}\cup((A\cap X^{\prime})\setminus V(Q_{1}\cup\ldots\cup Q_{t}))$. Together with Lemma 6.3 (iii), the desired rainbow directed path can be obtained immediately.
Finally, we consider $|\hat{X}|=0$ and $|B\cap X^{\prime}|\geq 1$. In this case, the desired rainbow directed path also follows immediately from part (iii) of Lemma 6.3. (Note that if there is a rainbow directed path with some endpoint in $V_{\rm bad}$, then it can be extended by at most two vertices so that both of its endpoints remain in the corresponding part and are no longer in $V_{\rm bad}$.) ∎
By Lemma 6.1 and a similar discussion as Steps 1-3 in Case 1, the following hold.
@@
    In $\{D_{i}[A\setminus V(P^{0})]\cup D_{i}[B\setminus V(P^{0})]:i\in\mathcal{C}_{1}\setminus{\rm col}(P^{0})\}$, there are $|V_{\rm bad}\setminus V(P^{0})|$ disjoint rainbow directed $P_{3}$ copies, each with its center in $V_{\rm bad}\setminus V(P^{0})$ and both endpoints in the same part. Let $\mathbf{P}$ be the set of these rainbow directed $P_{3}$ copies.
@@
    For colors in $\mathcal{C}_{\rm bad}$, there exists a maximal rainbow matching, say $M$, inside $\{D_{i}[A\setminus V(\mathbf{P}\cup P^{0})]\cup D_{i}[B\setminus V(\mathbf{P}\cup P^{0})]:i\in\mathcal{C}_{\rm bad}\}$ such that for each edge in $M[Y]$ with color $j$ we have $D_{j}[Y\setminus V(\mathbf{P}\cup P^{0}\cup{M})]$ contains a $2$-matching. Denote $\mathcal{C}_{\rm bad}^{\prime}:=\mathcal{C}_{\rm bad}\setminus{\rm col}(M)$. Then $D_{j}$ is $(19\sqrt{\delta},{\rm EC}2)$-extremal for all $j\in\mathcal{C}_{\rm bad}^{\prime}$.
@@
    Let $q:=|\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime}|$. In the digraph collection $\{D_{i}^{\pm}[A,B]:i\in\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime})\}$, there exists a rainbow directed path $Q$ of length q1q-1 or $q$ that avoids vertices in $\mathbf{P}\cup P^{0}\cup M$.
@@
Next, we give a claim that will be used later to obtain transversal directed Hamilton cycles.
@@
$\geq d_{i}^{+}(z,A_{i})-|A_{1}\triangle A_{i}|-|X_{A}\setminus X_{A}^{1}|-|X^{\prime}|-|V(Q_{1}\cup Q_{2})|-|I_{3-j}|$
@@
Similarly, we have $|N_{{i}}^{-}(z,I_{j})|\geq\frac{n}{4}-5\eta n$. This implies that
@@
Hence, for each color $i\in\mathcal{\mathcal{C}}_{a}\cup\{i_{1},i_{2}\}$, one may view $D_{i}^{\pm}[I_{1},I_{2}]$ as an undirected graph $G_{i}[I_{1},I_{2}]$ with minimum degree at least $\frac{n}{4}-10\eta n$, where two vertices are adjacent in $G_{i}[I_{1},I_{2}]$ if and only if there are two opposite arcs between them. By Lemma 6.2, there exists a transversal undirected path $Q_{a}$ inside $\mathcal{D}_{a}$ with two endpoints $x_{a}\in N_{G_{i_{1}}}(z_{1},I_{1})$ and $y_{a}\in N_{G_{i_{2}}}(z_{1}^{\prime},I)$ (here $y_{a}\in I_{2}$ if and only if $|I_{1}|=|I_{2}|$).
@@
Using Lemma 6.3 (iii), we connect $P^{0}$ (with endpoints in the same part), all rainbow directed paths in $\mathbf{P},$ all edges in $M$ and $Q$ (from $A$ to $B$ with length q1q-1) into a single rainbow directed path $P$, whose length is at most $(8\sqrt{\epsilon}+3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$ and endpoints are $x\in A\setminus(V_{\rm bad}\cup X^{\prime})$, $y\in B\setminus(V_{\rm bad}\cup X^{\prime})$. Choose $c\in(\mathcal{C}_{2}\cup\mathcal{C}_{\rm bad}^{\prime})\setminus{\rm col}(P)$ and let $u_{q}v_{q}$ be an edge in $D_{c}[B\setminus V(P),A\setminus V(P)]$. Consequently, $P$ and $u_{q}v_{q}$ are two disjoint rainbow directed paths satisfying all conditions in Claim 6.9. Thus, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
Using Lemma 6.3 (iii), we connect all rainbow paths in $\mathbf{P},$ all edges in $M$ and $Q$ (from $B$ to $A$ with length $q$) into a single rainbow directed path $P$, whose length is at most $(3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$ and endpoints are $x\in B\setminus V_{\rm bad}$, $y\in A\setminus V_{\rm bad}$. Therefore, $P$ and $P^{0}$ (from $A$ to $B$) are two disjoint rainbow paths satisfying all conditions in Claim 6.9. It follows that $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
If $|X^{\prime}|\geq 2$, then choose $u\in A\cap X^{\prime}$ if $|A\cap X^{\prime}|>|B\cap X^{\prime}|$, and choose $u\in B\cap X^{\prime}$ if $|A\cap X^{\prime}|=|B\cap X^{\prime}|$. Delete $u$ from $X^{\prime}$ when constructing $P^{0}$. Based on the definition of $X^{\prime}$, there exists a rainbow directed copy of $P_{3}$ from $B$ to $A$ with center $u$, whose vertices are disjoint from $P^{0}$. Applying Lemma 6.3 (iii) again to connect $P^{0}$ (from $A$ to $B$), all rainbow directed paths in $\mathbf{P},M$ into a single rainbow path $P$, whose length is at most $(8\sqrt{\epsilon}+3\sqrt{\delta}+16\sqrt{\delta}+\eta)n$, with endpoints in different parts. Thus, $P$ and $P_{x}$ are two disjoint rainbow paths satisfying all conditions in Claim 6.9. Therefore, $\mathcal{D}$ contains a transversal directed Hamilton cycle, a contradiction.
@@
Therefore, there exists a set of disjoint rainbow directed paths $\mathbf{P}^{0}=\{v_{i}^{1}v_{i}v_{i}^{2}:i\in[t]\}$ with ${\rm col}(v_{i}^{1}v_{i}v_{i}^{2})=\{c_{4i-3},c_{4i-2}\}$ for each $i\in[t]$.
@@
Based on the definition of $X^{\prime}$, we move a vertex from $X^{\prime}$ to $Y$ whenever it is covered by a rainbow directed $P_{3}$ from $Z$ to $Y$ (in Step 2 of Case 1). Avoiding all vertices and colors appearing in $\mathbf{P}^{0}$, we then obtain a family of disjoint rainbow directed paths $\mathbf{P^{1}}=\{v^{1}vv^{2}:v\in X^{\prime}\}$, where for each vertex $v\in Y\cap X^{\prime}$ we have $v^{1}\in Z\setminus(V_{\rm bad}\cup X^{\prime})$ and $v^{2}\in Y\setminus(V_{\rm bad}\cup X^{\prime})$. Furthermore, for every such path at least one of the colors of $v^{1}v$} and $vv^{2}$} can be chosen from both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. Without loss of generality, assume that $|A|\leq|B|$.
@@
Using colors in $\mathcal{C}_{1}\cup\mathcal{C}_{2}$ and applying Lemma 6.3 (i)-(iii), we connect all rainbow directed paths in $\mathbf{P}^{0}\cup\mathbf{P}^{1}$ into a single rainbow directed path $P^{1}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{1}\in B\setminus V_{\rm bad}$, whose length is at most $4t+4|X^{\prime}|\leq 2\sqrt{\delta}n+4|X^{\prime}|$.
@@
By Lemma 6.1, there exists a set of disjoint rainbow directed paths $\mathbf{P}^{2}=\{x_{i}^{1}x_{i}x_{i}^{2}:i\in[s]\}$, each of which has colors $c_{i}^{1},c_{i}^{2}$ for $i\in[s]$. Applying Lemma 6.3 (i)-(ii), we connect all these rainbow directed paths together with $P^{1}$ into a single rainbow directed path $P^{2}$, which has endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{2}\in B\setminus V_{\rm bad}$ and length at most $|E(P^{1})|+4|X\setminus X^{\prime}|+2\leq(2\sqrt{\delta}+4\sqrt{\epsilon})n+2$.
Choose a maximal rainbow matching, say $M$, inside $\{D_{i}[A\setminus V(P^{2})]\cup D_{i}[B\setminus V(P^{2})]:i\in\mathcal{C}_{\rm bad}\}$ such that for each edge in $M[Y]$ with color $j$, we have $D_{j}[Y\setminus V(P^{2}\cup M)]$ contains a $2$-matching. Denote $\mathcal{C}_{\rm bad}^{\prime}:=\mathcal{C}_{\rm bad}\setminus{\rm col}(M)$. Note that $D_{j}$ is $(19\sqrt{\delta},{\rm EC}2)$-extremal for all $j\in\mathcal{C}_{\rm bad}^{\prime}$. Hence, $\{D_{i}[A\setminus V(P^{2}\cup M),B\setminus V(P^{2}\cup M)]:i\in\mathcal{C}_{\rm bad}^{\prime}\}$ contains a transversal matching, say $M^{\prime}$. By using Lemma 6.3 (i)-(iii) to connect $P^{2}$ and all rainbow edges in $M\cup M^{\prime}$, we obtain a rainbow directed path $P^{3}$ with endpoints $u_{1}\in A\setminus V_{\rm bad}$ and $w_{3}\in B\setminus V_{\rm bad}$, whose length is at most $|E(P^{2})|+16\sqrt{\delta}n\leq(18\sqrt{\delta}+4\sqrt{\epsilon})n+2$.